OpenStudy (anonymous):

In right triangle DEF, d = 8, e = 10, f = 6. Which is cos D + cos F?

hero (hero):

Ah, this is different. I think D and F are angles to be found

hero (hero):

but which angle is the right triangle...I assume E

OpenStudy (anonymous):

I think so, there is no picture.

hero (hero):

\[D = \sin^{-1}(\frac {8}{10}) , F = \sin^{-1} (\frac {6}{10}) \]

hero (hero):

I get sin D + cos F = 1.6